# Designing Human-AI Systems: Anthropomorphism and Framing Bias on Human-AI Collaboration

### v1

@@ 4.3 Conclusions
All in all, this work points to the need to further investigate the role of cognitive biases in human-AI collaboration in order to design AI products that effectively interact with humans, while avoiding ethical pitfalls. The results suggest that, on certain occasions, humans are more likely to agree with the AI when it does not have an identity. In addition, subtle framing manipulations of AI outputs do not seem to significantly modify this agreement. However, the main limitations of this work, including its inability to replicate framing effects, the high rate of subjects failing attention checks, and the lack of studies in this area, indicate that there is still much to be discovered in the realm of human-AI collaboration. By continuing to explore this dynamic, we can develop AI systems that are better designed to work collaboratively with humans in a variety of contexts.
@@ Acknowledgements
The author thanks Dr. Michele Piazzai for constructive feedback on previous versions of this article and generous guidance.

### v2

@@ 4.4 Conclusions
All in all, this work points to the need to further investigate the role of cognitive biases in human-AI collaboration in order to design AI products that effectively interact with humans, while avoiding ethical pitfalls. The results suggest that, on certain occasions, humans are more likely to agree with the AI when its design lacks a specific set of human-like characteristics. In addition, subtle framing manipulations of AI outputs do not seem to significantly affect human decisions. However, the lack of studies in this area and the main limitations of this work, including its inability to replicate framing effects, the large number of subjects who failed the attention checks, and the lack of manipulation checks, indicate that there is still much to be discovered in the realm of human-AI collaboration. By continuing to explore this dynamic, we can develop AI systems that are better designed to work collaboratively with humans in a variety of contexts.
@@ Acknowledgements
We thank Michele Piazzai for constructive feedback on previous versions of this article and generous guidance.